c.)“Give me $10,000 today and I'll return $16,000 to you in five years," offers your investment broker. To the nearest tenth of a percent, what annual interest rate is being offered?

d.)How much money would you have to put away at the end of each year to have $1,600,000 when you retire 26 years from now if you can earn 4% on your money?

e.)How much can be accumulated if $2,375 is deposited at the end of each month for the next 28 years, if the account earns 4.2% APR (i.e. 0.35% per month).

f.)Assume the total expense for your current year in college equals $35,000. Approximately how much would you have needed to invest 8 years ago in an account paying 6.0% compounded annually to cover this amount?
Answer #1

c) The adjusted Annual return would be

16,000 (1/5)1 10,000

= 9.8560%

d) This can be done by the EMI formula

EMI = \frac{P * R * (1+R)^{N}}{((1+R)^{N} - 1)}

EMI = \frac{1600000 * 0.04 * (1+0.04)^{26}}{((1+0.04)^{26} - 1)}

= $100, 107.8087

e) Same formula now the EMI is given

EMI = \frac{P * R * (1+R)^{N}}{((1+R)^{N} - 1)}

2375 = \frac{P * 0.042 * (1+0.042)^{28}}{((1+0.042)^{28} - 1)}

P = $38,677.88

f) By compunding formula,

A = P * (1+R)^{N}

P = \frac{A}{(1+R)^{N}}

P = \frac{35000}{(1+0.06)^{8}}

= $ 21,959.433